# Differential Privacy of Cross-Attention with Provable Guarantee

### v1

@@ 6 Technique Overview
This section provides the technique overview of our paper. In Section 6.1, we analyze our DPTree data structure. In Section 6.2, we show the sensitivity of summation problem. In Section 6.3, we explain the high-level idea behind the weighted $\ell_{p}^{p}$ distance query. In Section 6.4, we show how to answer one-dimensional weighted $\ell_{1}$ distance query. In Section 6.5, we show how to answer softmax distance query using previous algorithms. In Section 6.6, we show how to handle adaptive query. By combining the results from these sections, we prove the main results in Section 4.
6.1 Summation Segment Tree
@@
First, in order to solve the weighted distance problem, we need to have a basic DP algorithm (Algorithm 1 and 2) that can answer simple summation queries. After analyzing its DP and error in Section B, we state the data structure theorem.
Theorem 6.1 (DPTree data structure, informal version of Theorem B.1).
@@
There is a data structure (see DPTree in Algorithm 1, 2) that uses $O(n)$ spaces to support the following operations:
@@
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
@@
During the design of the data structure, we found an interesting property based on the parallel composition rule of DP Fact 3.11. We will now state the lemma with its proof provided in Section B.
Lemma 6.2 (Weighted sum of disjoint interval queries, informal version of Lemma B.8).
@@
Then, we have $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ (see Alg. 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
@@
From Lemma 6.2, we can see that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries, as long as the sum of squared weights is finite.
6.2 Sensitivity for Range Summation Problem
Our DP summation tree data structure DPTree (Algorithm 1 and 2) requires sensitivity parameter $\Delta$. In this section, we show that for the summation problem, we have the sensitivity $\Delta=2R$ if the input $X\in[-R,R]^{n}$.
Lemma 6.3 (Sensitivity of summation).
@@ 6.2 Sensitivity for Range Summation Problem
Let’s say 2 neighboring dataset $X$ and $X^{\prime}$ differ in $x_{i}$ and $x^{\prime}_{i}$ for some $i$ in the array $X$. Then for a summation problem, i.e. $f(X):=\sum_{i=1}^{n}x_{i}$, we have
@@
6.3 Weighted $\ell_{p}^{p}$ Distance Problem
Figure 1: The visualization of how to build $c_{j}$ of rounded dataset $X\in[0,1]^{10}$ and compute the weighted $\ell_{1}$ distance. The number above each $x_{i}$ is $w_{i}$. See Algorithm 6 for details. Suppose $y=0$. Then $\sum_{i=1}^{n}w_{i}|y-x_{i}|=0.1\cdot 2.2+0.3\cdot 3.1+0.3\cdot(-2)+0.3\cdot(-3)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5+0.9\cdot(-1)+0.9\cdot 1=4.4$. And $\sum_{j=0}^{n}|k-j|c_{j}/n=0.1\cdot 2.2+0.3\cdot(-1.9)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5=4.4$. See details in Lemma C.1.
In this section, we introduce the intuition behind the method for handling the weighted $\ell_{p}^{p}$ distance problem. The formal lemmas and proofs can be found in Section C.
@@ 6.3 Weighted $\ell_{p}^{p}$ Distance Problem
Lemma 6.4 (Weighted $\ell_{p}^{p}$-distance high dimension, informal version of Lemma C.2).
@@
Remark 6.5.
In Lemma 6.4, we first round the dataset. This rounding simplifies the calculation by reducing the number of possible positions to consider, from real values in $[0,R]^{d}$ to the total $O(nd)$ spots. However, it also introduces an error $O(R/n)$ for one data point. Then, for one spot in the rounded dataset, we sum over the weights of that spot and multiply the corresponding distance raised to the power of $p$. Additionally, since we are dealing with $\ell_{p}^{p}$ distance, the rounding error is also raised to the power of $p$.
6.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
Based on previous discussions in Section 6.1 and 6.3, we can now describe our one-dimensional weighted $\ell_{1}$ distance data structure, DPTreeDistance, presented in Algorithm 6 and 7, which generalizes the results from [8]. Drawing from the intuition in Section 6.3, the initialization process is as follows: first, we round each data point in the dataset to the nearest multiple of a small interval and build an array that aggregates the corresponding weights. This array is then fed into our DPTree data structure in Algorithm 1. At query time, we query the DPTree to obtain the aggregated weights within a small interval and multiply these weights by the distance to the query point. Furthermore, we also introduce a relative error parameter $\alpha$ to reduce the total number of queries to $O(\log(n)/\alpha)$ instead of querying all $n$ positions. We also analyze the DP and the error bound; see details in Section D.
Theorem 6.6 (DPTreeDistance data structure, informal version of Theorem D.6).
There is a data structure DPTreeDistance (Algorithm 6,7) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
@@ 6.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
@@
6.5 Softmax Activation
We then describe how we extend the previous results to softmax activation, i.e. exponential inner product function (Definition 4.1). From [6], we know that softmax activation can be approximated by polynomial kernel function $P(\cdot)$ with a certain error. The following lemma shows that we can transform weighted softmax queries into polynomial kernels. More specifically, we have one term that computes the weighted $\ell_{2}^{2}$ distance, which is the place where we add DP noises. Because of the decomposability of the $\ell_{p}^{p}$ distance, i.e. $\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p}$, we can easily extend the results of Section 6.4 to handle the $\ell_{2}^{2}$ distance query. After that, we compute the term for the weighted $\ell_{2}^{2}$ norms of approximation kernel exactly. Summing all these terms, with a certain error, we can answer the softmax query. Related proofs can be found in Section G.
Lemma 6.7 (Weighted softmax approximation, informal version of Lemma G.6).
Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma G.5. Then, we can approximate the exponential inner product using the polynomial kernel function:
@@ 6.5 Softmax Activation
6.6 Adaptive Query
We introduce how we can modify our algorithm to solve the adaptive query problem using some tools in [75]. Our approach is based on proving that our algorithm can handle any query within the query space with a certain error. Since adaptive queries must lie within this space, our algorithm can effectively handle them. In Section 6.5, we demonstrate our algorithm’s capability to answer weighted softmax distance queries with constant probability. We then use the Chernoff bound to boost the constant probability of our algorithm to a high probability. Next, we apply the notion of an $\epsilon_{0}$-net to bound all query points within the net using the union bound. Finally, we bound all points in the query space by utilizing the Lipschitz property of the weighted softmax distance function and introducing an additive error. See the proofs in Sections F and G.
Lemma 6.8 (Adaptive softmax, informal version of Lemma G.10).
@@ 6.6 Adaptive Query
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
@@
7 Discussion
How do we extend to self-attention?
Why not add noise in some other places?
Where and how to add DP noises is an important problem to ask during the DP algorithm design. In this paper, we consider the problem of $\sum_{i=1}^{n}w_{i}\exp(\langle x_{i},y\rangle/d)$ where $y,x_{i}\in[0,R]^{d}$ and $w\in[-R_{w},R_{w}]^{n}$ (Definition 4.1). Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorihtm 1 and 2). From Lemma 6.3 and the way we initialize DPTree in Algorithm 6, we can see that the sensitivity $\Delta$ of this problem is $2R_{w}$.
A simple method for adding noise involves adding $n$ noises to a length $n$ array, with each item $w_{i}\exp(\langle x_{i},y\rangle/d)$ for $i\in[n]$. However, this approach increases the error by a factor of $n$ by basic composition (Fact 3.10) and also makes the model dependent on the number of queries. Besides, it only supports a single query and requires rebuilding the tree for each new query, rendering it impractical.
In contrast, our current noise-adding technique (Lines 11 and 17 of Algorithm 1) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$. This method also supports multiple queries, eliminating the need to rebuild the tree each time.
8 Conclusion
Acknowledgement
Appendix
Roadmap.
The appendix is organized as follows. In Section A, we give the preliminary of our paper. In Section B, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee. In Section C, we show how to solve weighted distance problem. In Section D, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{1}$ distance problem with DP and accuracy guarantee. In Section E, we present the analysis of our DPTreeHighDim (Algorithm 8) data structure, which can address the high-dimensional $\ell_{1}$ distance problem while ensuring differential privacy and accuracy guarantees. In Section F, we show how we can handle adaptive query. In Section G, we show how to extend our algorithm to softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 4).
Appendix A More Preliminary
In Section A.1, we give the probability tools we use in the paper. In Section A.2, we provide the facts we use.
A.1 Probability Tools
In this section, we give several probability lemmas.
Lemma A.1 (Markov’s inequality).
If $x$ is a nonnegative random variable and $t>0$, we have
$\Pr[x\geq t]\leq\frac{\operatorname*{{\mathbb{E}}}[x]}{t}.$
Lemma A.2 (Chernoff bound, [13]).
$\Pr[x\geq(1+\delta)\mu]\leq\exp(-\delta^{2}\mu/3),$
and for all $10<\delta<$ 1
$\Pr[x\leq(1-\delta)\mu]\leq\exp(-\delta^{2}\mu/2).$
Lemma A.3 (Chebyshev’s inequality).
$\Pr[|x-\mu|\geq k\sigma]\leq\frac{1}{k^{2}}.$
A.2 Facts
Fact A.4 (Upper bound of exponential, Fact C.9 in [41]).
For $a\in\mathbb{R}$, $b\in\mathbb{R}$, $a,b\leq R$, where $R\geq 0$, we have
$|\exp(a)-\exp(b)|\leq\exp(R)|a-b|.$
Appendix B DPTree Algorithm
In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 1 and 2). In Section B.1, we give the theorem (Theorem B.1) of our data structure that can answer summation problem. In Section B.2, we improve our data structure from constant probability to high probability by applying Chernoff bound. In Section B.3, we give the analysis. In Section B.4, we show some results of our data structure if the input queries are disjoint.
B.1 Single Data Structure
@@
Theorem B.1 (DPTree data structure, formal version of Theorem 6.1).
There is a data structure (see DPTree in Algorithm 1, 2) that uses $O(n)$ spaces to support the following operations:
@@
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
@@
The proofs follow from combining Lemma B.3 (running time of initialization), Lemma B.4 (running time of query), Lemma B.5 (DP of query), and Lemma B.6 (error of query) together. ∎
B.2 Boost the Constant Probability to High Probability
@@
1:datastructure DPTreeHighProb $\triangleright$ Theorem B.2
@@
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(\log(1/\delta_{\mathrm{fail}}))}:\textsc{DPTree}$triangleright Alg. 1
@@
Theorem B.2 (High-probability).
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
Proof.
Note that our data structure (Theorem B.1) succeeds with probability $0.99$. The success of the algorithm (Theorem B.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma A.2). By repeating the data structure $O(\log(1/\delta_{\mathrm{fail}}))$ times and taking the median of the outputs, we boost the success probability. The details are following.
To boost the success probability, we assume the query is repeated $l$ times. Let $i\in[l]$, and let $z_{i}$ denote the indicator random variable for the success of the $i$-th instance of the data structure for a single query. Let $z=\sum_{i=1}^{l}z_{i}$ be the total success times. Since $p=\Pr[z_{i}=1]=0.99$, we can have $\mu=\operatorname*{{\mathbb{E}}}[z]=\sum_{i=1}^{l}p=lp$. Note that $p=0.99$. By setting $\delta=0.1$ and using Chernoff bound from Lemma A.2, we can show
$\Pr[z>l/2]\geq 1-\exp(-\delta^{2}lp/2).$
To ensure that failure probability is $\delta_{\mathrm{fail}}$, we have
$\exp(-\delta^{2}lp/2)=\delta_{\mathrm{fail}}.$
We can make this hold by choosing $l=O(\log(1/\delta_{\mathrm{fail}}))$.
By the DP basic composition rule (Fact 3.10), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem B.1. ∎
B.3 Algorithm of Data Structure
In this section, we analyze the accuracy, DP, and runtime of Algorithm 1 and 2.

### v2

@@ Appendix C Proof Outline
This section provides the proof outline of our paper. In Section C.1, we analyze our DPTree data structure. In Section C.2, we show the sensitivity of summation problem. In Section C.3, we explain the high-level idea behind the weighted $\ell_{p}^{p}$ distance query. In Section C.4, we show how to answer one-dimensional weighted $\ell_{1}$ distance query. In Section C.5, we show how to answer Softmax distance query using previous algorithms. In Section C.6, we show how to handle adaptive query. By combining the results from these sections, we prove the main results in Section 4.
C.1 Summation Segment Tree
@@
First, in order to solve the weighted distance problem, we need to have a basic DP algorithm (Algorithm 2) that can answer simple summation queries. After analyzing its DP and error in Section D, we state the data structure theorem.
Theorem C.1 (DPTree data structure, informal version of Theorem D.1).
@@
There is a data structure (see DPTree in Algorithm 2) that uses $O(n)$ spaces to support the following operations:
@@
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
@@
During the design of the data structure, we found an interesting property based on the parallel composition rule of DP Fact B.7. We will now state the lemma, whose proof is provided in Section D.
Lemma C.2 (Weighted sum of disjoint interval queries, informal version of Lemma D.8).
@@
Then, we have Alg. 2 is $(\epsilon,\delta)$-DP and output $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ with the error upper bounded by
@@
From Lemma C.2, we can see that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries, as long as the sum of squared weights is finite.
C.2 Sensitivity for Range Summation Problem
Our DP summation tree data structure DPTree (Algorithm 2) requires sensitivity parameter $\Delta$. In this section, we show that for the summation problem, we have the sensitivity $\Delta=2R$ if the input $X\in[-R,R]^{n}$.
Lemma C.3 (Sensitivity of summation).
@@ C.2 Sensitivity for Range Summation Problem
Let’s say two neighboring datasets $X$ and $X^{\prime}$ differ in $x_{i}$ and $x^{\prime}_{i}$ for some $i$ in the array $X$. Then for a summation problem, i.e. $f(X):=\sum_{i=1}^{n}x_{i}$, we have
@@
C.3 Weighted $\ell_{p}^{p}$ Distance Problem
In this section, we introduce the intuition behind the method for handling the weighted $\ell_{p}^{p}$ distance problem. The formal lemmas and proofs can be found in Section E.
@@ C.3 Weighted $\ell_{p}^{p}$ Distance Problem
Lemma C.4 (Weighted $\ell_{p}^{p}$-distance high dimension, informal version of Lemma E.2).
@@
Remark C.5.
In Lemma C.4, we first round the dataset. This rounding simplifies the calculation by reducing the number of possible positions to consider, from real values in $[0,R]^{d}$ to the total $O(nd)$ spots. However, it also introduces an error $O(R/n)$ for one data point. Then, for one spot in the rounded dataset, we sum over the weights of that spot and multiply the corresponding distance raised to the power of $p$. Additionally, since we are dealing with $\ell_{p}^{p}$ distance, the rounding error is also raised to the power of $p$.
Figure 1: The visualization of how to build $c_{j}$ of rounded dataset $X\in[0,1]^{10}$ and compute the weighted $\ell_{1}$ distance. The number above each $x_{i}$ is $w_{i}$. See Algorithm 5 for details. Suppose $y=0$. Then $\sum_{i=1}^{n}w_{i}|y-x_{i}|=0.1\cdot 2.2+0.3\cdot 3.1+0.3\cdot(-2)+0.3\cdot(-3)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5+0.9\cdot(-1)+0.9\cdot 1=4.4$. And $\sum_{j=0}^{n}|k-j|c_{j}/n=0.1\cdot 2.2+0.3\cdot(-1.9)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5=4.4$. See details in Lemma E.1.
C.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
Based on previous discussions in Section C.1 and C.3, we can now describe our one-dimensional weighted $\ell_{1}$ distance data structure, DPTreeDistance, presented in Algorithm 5 and 6, which generalizes the results from BLM+ [10]. Drawing from the intuition in Section C.3, the initialization process is as follows: first, we round each data point in the dataset to the nearest multiple of a small interval and build an array that aggregates the corresponding weights. This array is then fed into our DPTree data structure in Algorithm 2. At query time, we query the DPTree to obtain the aggregated weights within a small interval and multiply these weights by the distance to the query point. Furthermore, we also introduce a relative error parameter $\alpha$ to reduce the total number of queries to $O(\log(n)/\alpha)$ instead of querying all $n$ positions. We also analyze the DP and the error bound; see details in Section F.
Theorem C.6 (DPTreeDistance data structure, informal version of Theorem F.6).
There is a data structure DPTreeDistance (Algorithm 5 and 6) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
@@ C.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
@@
C.5 Softmax Activation
We then describe how we extend the previous results to Softmax activation, i.e. exponential inner product function (Definition 4.1). From AS [6], we know that Softmax activation can be approximated by polynomial kernel function $P(\cdot)$ with a certain error. The following lemma shows that we can transform weighted Softmax queries into polynomial kernels. More specifically, we have one term that computes the weighted $\ell_{2}^{2}$ distance, which is the place where we add DP noises. Because of the decomposability of the $\ell_{p}^{p}$ distance, i.e. $\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p}$, we can easily extend the results of Section C.4 to handle the $\ell_{2}^{2}$ distance query. After that, we compute the term for the weighted $\ell_{2}^{2}$ norms of approximation kernel exactly. Summing all these terms, with a certain error, we can answer the Softmax query. Related proofs can be found in Section I.
Lemma C.7 (Weighted Softmax approximation, informal version of Lemma I.6).
Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma I.5. Then, we can approximate the exponential inner product using the polynomial kernel function:
@@ C.5 Softmax Activation
C.6 Adaptive Query
We introduce how we can modify our algorithm to solve the adaptive query problem using some tools in QRS+ [86]. Our approach is based on proving that our algorithm can handle any query within the query space with a certain error. Since adaptive queries must lie within this space, our algorithm can effectively handle them. In Section C.5, we demonstrate our algorithm’s capability to answer weighted Softmax distance queries with constant probability. We then use the Chernoff bound to boost the constant probability of our algorithm to a high probability. Next, we apply the notion of an $\epsilon_{0}$-net to bound all query points within the net using the union bound. Finally, we bound all points in the query space by utilizing the Lipschitz property of the weighted Softmax distance function and introducing an additive error. See the proofs in Sections H and I.
Lemma C.8 (Adaptive Softmax, informal version of Lemma I.10).
@@ C.6 Adaptive Query
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
@@
Appendix D DPTree Algorithm
In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 2). In Section D.1, we give the theorem (Theorem D.1) of our data structure that can answer summation problem. In Section D.2, we improve our data structure from constant probability to high probability by applying Chernoff bound. In Section D.3, we give the analysis. In Section D.4, we show some results of our data structure if the input queries are disjoint.
D.1 Single Data Structure
@@
Theorem D.1 (DPTree data structure, formal version of Theorem C.1).
There is a data structure (see DPTree in Algorithm 2) that uses $O(n)$ spaces to support the following operations:
@@
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
@@
D.2 Boost the Constant Probability to High Probability
Theorem D.2 (High-probability).
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
Proof.
Note that our data structure (Theorem D.1) succeeds with probability $0.99$. The success of the algorithm (Theorem D.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma B.2). By repeating the data structure $O(\log(1/\delta_{\mathrm{fail}}))$ times and taking the median of the outputs, we boost the success probability. The details are following.
To boost the success probability, we assume the query is repeated $l$ times. Let $i\in[l]$, and let $z_{i}$ denote the indicator random variable for the success of the $i$-th instance of the data structure for a single query. Let $z=\sum_{i=1}^{l}z_{i}$ be the total success times. Since $p=\Pr[z_{i}=1]=0.99$, we can have $\mu=\operatorname*{{\mathbb{E}}}[z]=\sum_{i=1}^{l}p=lp$. Note that $p=0.99$. By setting $\delta=0.1$ and using Chernoff bound from Lemma B.2, we can show
$\Pr[z>l/2]\geq 1-\exp(-\delta^{2}lp/2).$
To ensure that failure probability is $\delta_{\mathrm{fail}}$, we have
$\exp(-\delta^{2}lp/2)=\delta_{\mathrm{fail}}.$
We can make this hold by choosing $l=O(\log(1/\delta_{\mathrm{fail}}))$.
By the DP basic composition rule (Fact B.6), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem D.1. ∎
@@ D.2 Boost the Constant Probability to High Probability
1:datastructure DPTreeHighProb $\triangleright$ Theorem D.2
@@
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(\log(1/\delta_{\mathrm{fail}}))}:\textsc{DPTree}$triangleright Alg. 2
@@
D.3 Algorithm of Data Structure
In this section, we analyze the accuracy, DP, and runtime of Algorithm 2.